# Customized SAM 2 for Referring Remote Sensing Image Segmentation

### v2

@@ 4.2 Implementation Details
Model Settings. We initialize the key modules of SAM 2 and the union encoder using pre-trained weights from SAM 2-Hiera-Large [24] and BEiT-3-Large [29]. For feature extraction, each image is resized to $1024\times 1024$ and $224\times 224$, which are then fed into the SAM 2 image encoder with an output dimension of $C=256$ and the union encoder with an output dimension of $D=1024$. Unlike SAM 2’s full capabilities, we do not employ its memory mechanism, focusing exclusively on image processing. In the bidirectional hierarchical fusion module, we set the text weight coefficient $\alpha_{t}$ to 0.2 and the image weight coefficient $\alpha_{i}$ to 0.5.
@@
Training Details. Experiments are conducted on 8 NVIDIA GeForce RTX 4090 GPUs. The experiment adopts a setup similar to [14], with training conducted for 60 epochs on the RefSegRS [38] dataset and 40 epochs on the RRSIS-D [19] dataset. We use the AdamW optimizer [21] with a unified batch size of 1. The learning rates for models on RefSegBS and RRSIS-D dataset are set to 5e-5 and 1e-5, respectively. To facilitate the adaptation of our designed modules to remote sensing data, the learning rates for bidirectional hierarchical fusion module and mask prompt generator are assigned as 1e-4 on RefSegRS dataset and 5e-5 on RRSIS-D dataset. All learning rates are reduced to 0.1 times their original values during the last 10 epochs. The weighting coefficients for various loss functions are defined as follows: $\lambda_{dice}=0.1$, $\lambda_{ce}=1$, and $\lambda_{tbl}=0.2$.
@@
Method Reference Pr@0.5 Pr@0.6 Pr@0.7 Pr@0.8 Pr@0.9 oIoU mIoU
@@
BRINet [6] CVPR’20 36.86 20.72 35.53 14.26 19.93 9.87 10.66 2.98 2.84 1.14 61.59 58.22 38.73 31.51
LSCM [8] ECCV’20 56.82 31.54 41.24 20.41 21.85 9.51 12.11 5.29 2.51 0.84 62.82 61.27 40.59 35.54
CMPC [7] CVPR’20 46.09 32.36 26.45 14.14 12.76 6.55 7.42 1.76 1.39 0.22 63.55 55.39 42.08 40.63
CMSA [36] CVPR’19 39.24 28.07 38.44 20.25 20.39 12.71 11.79 5.61 1.52 0.83 65.84 64.53 43.62 41.47
RRN [15] CVPR’18 55.43 30.26 42.98 23.01 23.11 14.87 13.72 7.17 2.64 0.98 69.24 65.06 50.81 41.88
EVF-SAM [43] Arxiv’24 57.77 35.17 37.59 22.34 16.24 9.36 4.87 2.86 1.86 0.39 59.61 55.51 46.98 36.64
CMPC+ [18] TPAMI’21 56.84 49.19 37.59 28.31 20.42 15.31 10.67 8.12 2.78 2.55 70.62 66.53 47.13 43.65
CARIS [20] ACM MM’23 68.45 45.40 47.10 27.19 25.52 15.08 14.62 8.87 3.71 1.98 75.79 69.74 54.30 42.66
CRIS [30] CVPR’22 53.13 35.77 36.19 24.11 24.36 14.36 11.83 6.38 2.55 1.21 72.14 65.87 53.74 43.26
LAVT [35] CVPR’22 80.97 51.84 58.70 30.27 31.09 17.34 15.55 9.52 4.64 2.09 78.50 71.86 61.53 47.40
RIS-DMMI [5] CVPR’23 86.17 63.89 74.71 44.30 38.05 19.81 18.10 6.49 3.25 1.00 74.02 68.58 65.72 52.15
LGCE [38] TGRS’24 90.72 73.75 86.31 61.14 71.93 39.46 32.95 16.02 10.21 5.45 83.56 76.81 72.51 59.96
RMSIN [19] CVPR’24 93.97 79.20 89.33 65.99 74.25 42.98 29.70 16.51 7.89 3.25 82.41 75.72 73.84 62.58
FIANet [14] TGRS’24 - 84.09 - 77.05 - 61.86 - 33.41 - 7.10 - 78.32 - 68.67
RS2-SAM 2 - 95.36 84.31 94.90 79.42 92.58 70.89 83.76 55.70 36.66 21.19 88.03 80.87 85.21 73.90
@@
Method Reference Pr@0.5 Pr@0.6 Pr@0.7 Pr@0.8 Pr@0.9 oIoU mIoU
Val Test Val Test Val Test Val Test Val Test Val Test Val Test
RRN [15] CVPR’18 51.09 51.07 42.47 42.11 33.04 32.77 20.80 21.57 6.14 6.37 66.53 66.43 46.06 45.64
CMSA [36] CVPR’19 55.68 55.32 48.04 46.45 38.27 37.43 26.55 25.39 9.02 8.15 69.68 69.39 48.85 48.54
LSCM [8] ECCV’20 57.12 56.02 48.04 46.25 37.87 37.70 26.37 25.28 7.93 8.27 69.28 69.05 50.36 49.92
CMPC [7] CVPR’20 57.93 55.83 48.85 47.40 38.50 36.94 25.28 25.45 9.31 9.19 70.15 69.22 50.41 49.24
BRINet [6] CVPR’20 58.79 56.90 49.54 48.77 39.65 39.12 28.21 27.03 9.19 8.73 70.73 69.88 51.14 49.65
CMPC$+$ [18] TPAMI’21 59.19 57.65 49.36 47.51 38.67 36.97 25.91 24.33 8.16 7.78 70.14 68.64 51.41 50.24
LGCE [38] TGRS’24 68.10 67.65 60.52 61.53 52.24 51.45 42.24 39.62 23.85 23.33 76.68 76.34 60.16 59.37
RIS-DMMI [5] CVPR’23 70.40 68.74 63.05 60.96 54.14 50.33 41.95 38.38 23.85 21.63 77.01 76.20 60.72 60.12
LAVT [35] CVPR’22 69.54 69.52 63.51 63.63 53.16 53.29 43.97 41.60 24.25 24.94 77.59 77.19 61.46 61.04
EVF-SAM [43] Arxiv’24 73.51 72.16 67.87 66.50 58.33 56.59 46.15 43.92 25.92 25.48 76.32 76.77 64.03 62.75
FIANet [14] TGRS’24 - 74.46 - 66.96 - 56.31 - 42.83 - 24.13 - 76.91 - 64.01
RMSIN [19] CVPR’24 74.66 74.26 68.22 67.25 57.41 55.93 45.29 42.55 24.43 24.53 78.27 77.79 65.10 64.20
RS2-SAM 2 - 79.25 77.56 74.08 72.34 63.85 61.76 50.57 47.92 30.40 29.73 80.16 78.99 68.81 66.72
4.3 Comparison with State-of-the-Art Methods
@@
RefSegRS dataset. We conduct a comprehensive comparison between our method, RS2-SAM 2, and several state-of-the-art approaches on the RefSegRS [38] dataset. The results of the comparison are reported in table 1. It can be observed that our RS2-SAM 2 achieves 88.03% oIoU and 85.21% mIoU on the validation set, surpassing the previous best method, RMSIN, by 5.62% in oIoU and 11.37% in mIoU. On the test set, it reaches 80.87% oIoU and 73.90% mIoU, outperforming FIANet [14] by 2.55% and 5.23%, respectively. Moreover, the Pr metric exhibits significant improvements across all thresholds, particularly at Pr@0.7, Pr@0.8, and Pr@0.9, demonstrating the model’s strong multimodal segmentation capabilities.
@@
RRSIS-D dataset. We also conduct comparative experiments between our RS2-SAM 2 and existing methods such as LGCE [38], LAVT [35], EVF-SAM [43] and RMSIN [19], etc., on the RRSIS-D [19] dataset, with results documented in table 2. On this dataset, our method achieves the best performance, surpassing the current state-of-the-art method RMSIN on the validation set by 4.59% in Pr@0.5, 5.86% in Pr@0.6, 6.44% in Pr@0.7, 5.28% in Pr@0.8, 5.97% in Pr@0.9, 1.89% in oIoU, and 3.71% in mIoU. It also significantly outperformed existing methods on the test set, demonstrating its strong capability in remote sensing object segmentation.
@@
fig. 4 illustrates the visual comparison between our model and RMSIN [19] on the RRSIS-D dataset. The results clearly indicate that RS2-SAM 2 consistently outperforms RMSIN, particularly in terms of accurate target localization, reliable mask prediction, and boundary precision.
Method $\mathcal{L}_{tbl}$ MMP DHF Pr@0.5 Pr@0.7 Pr@0.9 mIoU oIoU
@@
4.4 Model Analysis
In this section, we perform extensive ablation studies to analyze the impact of the essential components in our RS2-SAM 2 framework, as well as the effects of different model configurations. The experiments are conducted on the test dataset of RefSegRS [38].
Components Analysis. To investigate the impact of key components in our model, we first construct a baseline model consisting solely of SAM 2 and the union encoder. As shown in table 3, when the text-guided boundary loss ($\mathcal{L}_{tbl}$) is incorporated during training, the model achieves 38.63% mIoU and 57.36% oIoU, surpassing the baseline by 1.99% and 1.85%, respectively. Subsequently, on top of the model with $\mathcal{L}_{tbl}$, we add the mask prompt generator (MPG), which results in the RS2-SAM 2 achieving 60.20% mIoU and 70.89% oIoU, an improvement of 21.57% and 13.53% over the previous model. When the bidirectional hierarchical fusion module (BHFM) is added to the baseline model with $\mathcal{L}_{tbl}$, the model shows a further increase of 30.08% in mIoU and 21.00% in oIoU, further validating the superiority of this module. Finally, when all components are integrated, our RS2-SAM 2 achieves the best performance, with a mIoU of 73.90% and an oIoU of 80.87%.
Figure 4: Visualization result on RRSIS-D. Compared to RMSIN [19], RS2-SAM 2 demonstrates superior capability in handling local details and boundary regions.
Mask Prompt Generator. In this section, we explore different interaction forms within the mask prompt generator (MPG), with results presented in table 4. When the multi-head cross-attention (MHCA) between the multimodal token $V_{cls}$ and visual embeddings $V$ is omitted, the performance of RS2-SAM 2 drops by 2.31% in mIoU and 0.98% in oIoU. These findings underscore the importance of further strengthening the semantic connections between the multimodal token and visual embeddings during the multimodal mask prompt generation process.
Table 4: Model analysis of different settings in RS2-SAM 2.
@@ 4.4 Model Analysis
Bidirectional Hierarchical Fusion Module. The impact of different settings of the bidirectional hierarchical fusion module (BHFM) is also noteworthy. First, we examine how the structure of the BHFM layer during feature encoding affects the model’s performance. One simple approach is to use a linear layer structure similar to the adapter, without interacting with text feature, referred to as “Linear". Another approach is to enhance visual feature using only text feature, without back-propagating visual information, referred to as “Uni". The third approach is bidirectional enhancement of both text and visual features, referred to as “Bi". The experimental results in table 4 show that the layer-by-layer bidirectional enhancement of text and visual features is more beneficial for imparting linguistic information to the visual feature. This enables refined linguistic enhancement of the visual feature from low-level to high-level, leading to more accurate segmentation of remote sensing objects.
Additionally, we investigate the impact of BHFM components on segmentation performance by setting up experiments where the BHFM cross-attention (BC) after encoding and the BHFM layer (BL) without the encoder are removed. Results in table 4 indicate that both interactions during and after encoding are essential. The hierarchical interaction helps the model understand text features from a global to local perspective, enhancing its pixel-level understanding of text.
5 Conclusion
In this paper, we present RS2-SAM 2, an advanced end-to-end framework designed to enhance SAM 2 for RRSIS. Our approach leverages a union encoder to jointly encode visual and textual features, producing semantically aligned visual-text embeddings and multimodal class tokens. To effectively integrate spatial and textual information, we introduce a bidirectional hierarchical fusion module, which incorporates textual semantics and spatial context both during and after encoding, enabling a hierarchical refinement from global to local levels. Additionally, a mask prompt generator generates multimodal mask as dense prompt, improving the segmentation of visually indistinct objects by providing stronger pixel-level guidance. To further enhance boundary precision, we propose a text-guided boundary loss, which refines segmentation results by computing text-weighted boundary gradient differences. Extensive experiments on multiple RRSIS benchmarks demonstrate the superiority of RS2-SAM 2 over state-of-the-art methods, validating the effectiveness of our proposed modules.

### v3

@@ Introduction
Model Settings. We initialize the key modules of SAM2 and the union encoder using pre-trained weights from SAM2-Hiera-Large (Ravi et al. 2024) and BEiT-3-Large (Wang et al. 2023). For feature extraction, each image is resized to $\times 1024$ 1024 and $\times 224$ 224, which are then fed into the SAM2 image encoder with an output dimension of $C=256$ and the union encoder with an output dimension of $D=1024$. Unlike SAM2’s full capabilities, we do not employ its memory mechanism, focusing exclusively on image processing. In the bidirectional hierarchical fusion module, we set the text weight coefficient $\alpha_{t}$ to 0.2 and the image weight coefficient $\alpha_{i}$ to 0.5.
@@
Training Details. Experiments are conducted on 8 NVIDIA GeForce RTX 4090 GPUs. The experiment adopts a setup similar to (Lei et al. 2024), with training conducted for 60 epochs on the RefSegRS (Yuan et al. 2024) dataset and 40 epochs on the RRSIS-D (Liu et al. 2024b) dataset. We use the AdamW optimizer (Loshchilov and Hutter 2018) with a unified batch size of 1. The learning rates for models on RefSegBS and RRSIS-D dataset are set to 5e-5 and 1e-5, respectively. To facilitate the adaptation of our designed modules to remote sensing data, the learning rates for bidirectional hierarchical fusion module and mask prompt generator are assigned as 1e-4 on RefSegRS dataset and 5e-5 on RRSIS-D dataset. All learning rates are reduced to 0.1 times their original values during the last 10 epochs. The weighting coefficients for various loss functions are defined as follows: $\lambda_{dice}=0.1$, $\lambda_{ce}=1$, and $\lambda_{tbl}=0.2$.
Method Pr@0.5 Pr@0.6 Pr@0.7 Pr@0.8 Pr@0.9 oIoU mIoU
Val Test Val Test Val Test Val Test Val Test Val Test Val Test
BRINet (Hu et al. 2020) 36.86 20.72 35.53 14.26 19.93 9.87 10.66 2.98 2.84 1.14 61.59 58.22 38.73 31.51
LSCM (Hui et al. 2020) 56.82 31.54 41.24 20.41 21.85 9.51 12.11 5.29 2.51 0.84 62.82 61.27 40.59 35.54
CMPC (Huang et al. 2020) 46.09 32.36 26.45 14.14 12.76 6.55 7.42 1.76 1.39 0.22 63.55 55.39 42.08 40.63
CMSA (Ye et al. 2019) 39.24 28.07 38.44 20.25 20.39 12.71 11.79 5.61 1.52 0.83 65.84 64.53 43.62 41.47
RRN (Li et al. 2018) 55.43 30.26 42.98 23.01 23.11 14.87 13.72 7.17 2.64 0.98 69.24 65.06 50.81 41.88
EVF-SAM (Zhang et al. 2024) 57.77 35.17 37.59 22.34 16.24 9.36 4.87 2.86 1.86 0.39 59.61 55.51 46.98 36.64
CMPC+ (Liu et al. 2021) 56.84 49.19 37.59 28.31 20.42 15.31 10.67 8.12 2.78 2.55 70.62 66.53 47.13 43.65
CARIS (Liu et al. 2023b) 68.45 45.40 47.10 27.19 25.52 15.08 14.62 8.87 3.71 1.98 75.79 69.74 54.30 42.66
CRIS (Wang et al. 2022) 53.13 35.77 36.19 24.11 24.36 14.36 11.83 6.38 2.55 1.21 72.14 65.87 53.74 43.26
LAVT (Yang et al. 2022) 80.97 51.84 58.70 30.27 31.09 17.34 15.55 9.52 4.64 2.09 78.50 71.86 61.53 47.40
RIS-DMMI (Hu et al. 2023) 86.17 63.89 74.71 44.30 38.05 19.81 18.10 6.49 3.25 1.00 74.02 68.58 65.72 52.15
LGCE (Yuan et al. 2024) 90.72 73.75 86.31 61.14 71.93 39.46 32.95 16.02 10.21 5.45 83.56 76.81 72.51 59.96
RMSIN (Liu et al. 2024b) 93.97 79.20 89.33 65.99 74.25 42.98 29.70 16.51 7.89 3.25 82.41 75.72 73.84 62.58
FIANet (Lei et al. 2024) - 84.09 - 77.05 - 61.86 - 33.41 - 7.10 - 78.32 - 68.67
RS2-SAM2 95.36 84.31 94.90 79.42 92.58 70.89 83.76 55.70 36.66 21.19 88.03 80.87 85.21 73.90
@@
Method Pr@0.5 Pr@0.6 Pr@0.7 Pr@0.8 Pr@0.9 oIoU mIoU
@@
RRN (Li et al. 2018) 51.09 51.07 42.47 42.11 33.04 32.77 20.80 21.57 6.14 6.37 66.53 66.43 46.06 45.64
CMSA (Ye et al. 2019) 55.68 55.32 48.04 46.45 38.27 37.43 26.55 25.39 9.02 8.15 69.68 69.39 48.85 48.54
LSCM (Hui et al. 2020) 57.12 56.02 48.04 46.25 37.87 37.70 26.37 25.28 7.93 8.27 69.28 69.05 50.36 49.92
CMPC (Huang et al. 2020) 57.93 55.83 48.85 47.40 38.50 36.94 25.28 25.45 9.31 9.19 70.15 69.22 50.41 49.24
BRINet (Hu et al. 2020) 58.79 56.90 49.54 48.77 39.65 39.12 28.21 27.03 9.19 8.73 70.73 69.88 51.14 49.65
CMPC$+$ (Liu et al. 2021) 59.19 57.65 49.36 47.51 38.67 36.97 25.91 24.33 8.16 7.78 70.14 68.64 51.41 50.24
LGCE (Yuan et al. 2024) 68.10 67.65 60.52 61.53 52.24 51.45 42.24 39.62 23.85 23.33 76.68 76.34 60.16 59.37
RIS-DMMI (Hu et al. 2023) 70.40 68.74 63.05 60.96 54.14 50.33 41.95 38.38 23.85 21.63 77.01 76.20 60.72 60.12
LAVT (Yang et al. 2022) 69.54 69.52 63.51 63.63 53.16 53.29 43.97 41.60 24.25 24.94 77.59 77.19 61.46 61.04
EVF-SAM (Zhang et al. 2024) 73.51 72.16 67.87 66.50 58.33 56.59 46.15 43.92 25.92 25.48 76.32 76.77 64.03 62.75
FIANet (Lei et al. 2024) - 74.46 - 66.96 - 56.31 - 42.83 - 24.13 - 76.91 - 64.01
RMSIN (Liu et al. 2024b) 74.66 74.26 68.22 67.25 57.41 55.93 45.29 42.55 24.43 24.53 78.27 77.79 65.10 64.20
RS2-SAM2 79.25 77.56 74.08 72.34 63.85 61.76 50.57 47.92 30.40 29.73 80.16 78.99 68.81 66.72
@@
Comparison with State-of-the-Art Methods
@@
RefSegRS dataset. We conduct a comprehensive comparison between our method, RS2-SAM2, and several state-of-the-art approaches on the RefSegRS (Yuan et al. 2024) dataset. The results of the comparison are reported in Tab. 1. It can be observed that our RS2-SAM2 achieves 88.03% oIoU and 85.21% mIoU on the validation set, surpassing the previous best method, RMSIN, by 5.62% in oIoU and 11.37% in mIoU. On the test set, it reaches 80.87% oIoU and 73.90% mIoU, outperforming FIANet (Lei et al. 2024) by 2.55% and 5.23%, respectively. Moreover, the Pr metric exhibits significant improvements across all thresholds, particularly at Pr@0.7, Pr@0.8, and Pr@0.9, demonstrating the model’s strong multimodal segmentation capabilities.
@@
RRSIS-D dataset. We also conduct comparative experiments between our RS2-SAM2 and existing methods such as LGCE (Yuan et al. 2024), LAVT (Yang et al. 2022), EVF-SAM (Zhang et al. 2024) and RMSIN (Liu et al. 2024b), etc., on the RRSIS-D (Liu et al. 2024b) dataset, with results documented in Tab. 2. On this dataset, our method achieves the best performance, surpassing the current state-of-the-art method RMSIN on the validation set by 4.59% in Pr@0.5, 5.86% in Pr@0.6, 6.44% in Pr@0.7, 5.28% in Pr@0.8, 5.97% in Pr@0.9, 1.89% in oIoU, and 3.71% in mIoU. It also significantly outperformed existing methods on the test set, demonstrating its strong capability in remote sensing object segmentation.
@@
Fig. 4 illustrates the visual comparison between our model and RMSIN (Liu et al. 2024b) on the RRSIS-D dataset. The results clearly indicate that RS2-SAM2 consistently outperforms RMSIN, particularly in terms of accurate target localization, reliable mask prediction, and boundary precision.
Figure 4: Visualization result on RRSIS-D. Compared to RMSIN (Liu et al. 2024b), RS2-SAM2 demonstrates superior capability in handling local details and boundary regions.
Model Analysis
In this section, we perform extensive ablation studies to analyze the impact of the essential components in our RS2-SAM2 framework, as well as the effects of different model configurations. The experiments are conducted on the test dataset of RefSegRS (Yuan et al. 2024).
Components Analysis. To investigate the impact of key components in our model, we first construct a baseline model consisting solely of SAM2 and the union encoder. As shown in Tab. 3, when the text-guided boundary loss ($\mathcal{L}_{tbl}$) is incorporated during training, the model achieves 38.63% mIoU and 57.36% oIoU, surpassing the baseline by 1.99% and 1.85%, respectively. Subsequently, on top of the model with $\mathcal{L}_{tbl}$, we add the mask prompt generator (MPG), which results in the RS2-SAM2 achieving 60.20% mIoU and 70.89% oIoU, an improvement of 21.57% and 13.53% over the previous model. When the bidirectional hierarchical fusion module (BHFM) is added to the baseline model with $\mathcal{L}_{tbl}$, the model shows a further increase of 30.08% in mIoU and 21.00% in oIoU, further validating the superiority of this module. Finally, when all components are integrated, our RS2-SAM2 achieves the best performance, with a mIoU of 73.90% and an oIoU of 80.87%.
Method $\mathcal{L}_{tbl}$ MPG BHFM Pr@0.5 Pr@0.7 Pr@0.9 mIoU oIoU
@@
Mask Prompt Generator. In this section, we explore different interaction forms within the mask prompt generator (MPG), with results presented in Tab. 4. When the multi-head cross-attention (MHCA) between the multimodal token $V_{cls}$ and visual embeddings $V$ is omitted, the performance of RS2-SAM2 drops by 2.31% in mIoU and 0.98% in oIoU. These findings underscore the importance of further strengthening the semantic connections between the multimodal token and visual embeddings during the multimodal mask prompt generation process.
Bidirectional Hierarchical Fusion Module. The effect of different BHFM configurations is also investigated. We first assess how the BHFM structure in the encoder influences performance. A simple variant uses a linear adapter-like layer without text interaction (“Linear”). Another uses text to enhance visual features without feedback (“Uni”). The third performs bidirectional enhancement between text and vision (“Bi”). As shown in Tab. 4, the layer-by-layer bidirectional enhancement proves more effective in infusing linguistic cues into visual features, enabling progressive textual refinement from low to high levels and resulting in more precise remote sensing segmentation.
@@
Table 4: Model analysis of different settings in RS2-SAM2.
Additionally, we investigate the impact of BHFM components on segmentation performance by setting up experiments where the BHFM cross-attention (BC) after encoding and the BHFM layer (BL) without the encoder are removed. The results in Tab. 4 indicate that both interactions during and after encoding are essential. The hierarchical interaction combining both can help the model understand text features from a global to a local perspective, enhancing its pixel-level understanding of text.
Conclusion
In this paper, we present RS2-SAM2, an advanced end-to-end framework designed to enhance SAM2 for RRSIS. Our approach leverages a union encoder to jointly encode visual and textual features, producing semantically aligned visual-text embeddings and multimodal class tokens. To effectively integrate spatial and textual information, we introduce a bidirectional hierarchical fusion module, which incorporates textual semantics and spatial context both during and after encoding, enabling a hierarchical refinement from global to local levels. Additionally, a mask prompt generator generates multimodal mask as dense prompt, improving the segmentation of visually indistinct objects by providing stronger pixel-level guidance. Extensive experiments on multiple RRSIS benchmarks demonstrate the superiority of RS2-SAM2 over state-of-the-art methods, validating the effectiveness of our proposed modules.